OpenStudy (anonymous):

Rewrite the following differential equation as an equivalent system of first-order differential equations. Use the variables x1=x x2=x' x3=x'', etc. (so that your equations will have the form x2'=x3, etc.: denote subscripts by appending the subscript to the variable name: X1= x1) Equation to rewrite: x''''+20x''−13x'+17x=−5cos(9t) x1' = x2' = x3' = x4' =

OpenStudy (eyust707):

\(x _{1} = x\) \(x _{2} = x'\) \(x _{3} = x''\) \(x _{4} = x'''\) \(x _{5} = x''''\)

OpenStudy (anonymous):

how do solve for those values though

OpenStudy (eyust707):

Im not exactly sure... I know they are saying that \[x _{2}'=x'' = x _{3}\]

OpenStudy (anonymous):

yea i got that too but im not sure what else were supposed to be doing ijn order to get those values

OpenStudy (eyust707):

im gunna go read a little and try and figure this out. maybe by that time someone like @TuringTest or @dpaInc can help you

OpenStudy (anonymous):

ok thank you and I will continue reading as well

OpenStudy (eyust707):

what is the section of your diffs book that this cam from titled?

OpenStudy (anonymous):

it wasn't from my book it was from an online web homework site and ive been looking in the book for similar problems and can't find it

OpenStudy (anonymous):

but i assume it has something to do with first-order systems

OpenStudy (eyust707):

is this the first one in the homework?

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

Simply rearranging we can say this: \(x''''=-20x''+13x'-17x−5\cos(9t)\) Then using what we said earlier: \(x_{4}' = -20x''+13x'-17x−5\cos(9t)\)

OpenStudy (eyust707):

But now we are second order

OpenStudy (anonymous):

i do not think thats what its asking. I do believe it is first order systems

OpenStudy (eyust707):

so: \[x_{4}' = -20x_{2}'+13x'-17x−5\cos(9t)\]

OpenStudy (eyust707):

and your system also includes these equations:

OpenStudy (anonymous):

yes but i think what we're supposed to it is know that x''''=f(t,x,x',x'',x''') and then we use what they gave us to solve for the answers. we know that x1=x and x2=x'=x1' and x3=x''=x2' and x4=x''''=x1'''=x2''=x3'

OpenStudy (anonymous):

so we get x1'=x2 x2'=x3 x3'=x4 x4'=?

OpenStudy (eyust707):

\(x_{1}' = x_{2}\) \(x_{2}' = x_{3}\) \(x_{3}' = x_{4}\) \(x_{4}' =-20x_{2}'+13x'-17x−5\cos(9t) \)

OpenStudy (anonymous):

do we need to substitute anything in

OpenStudy (eyust707):

honestly im not exactly sure what form they want the variables of the last equation

OpenStudy (eyust707):

We are definatley close tho... lol

OpenStudy (anonymous):

lol yes we are thank you

OpenStudy (eyust707):

In this one they put them in terms of \(x_{1}\) and \(x_{2}\) http://www.phaser.com/modules/elaydi/content/helptips/secondorderode.html

OpenStudy (eyust707):

plus all the other equations are in those terms... so might as well.

OpenStudy (anonymous):

ok let me try it and see what happens

OpenStudy (anonymous):

woo it works!! I guess it wasnt as complicated as it seemed to be! we didhave to replace the variables

OpenStudy (eyust707):

=P

OpenStudy (eyust707):

yes diffs is weird stuff... just flat out weird..

OpenStudy (anonymous):

yea i agree! lol

OpenStudy (eyust707):

whats even weirder is that theres all sorts of natural phenomenon that somehow fit themselves into the solution to differential equations...

OpenStudy (anonymous):

i know!

OpenStudy (anonymous):

differential equations is a very interesting subject